OpenStudy (anonymous):

Given the function g(x) = −2x + 4, find g(−5)

OpenStudy (anonymous):

g(-5)= -2(-5)+4 = 10+4= 14

OpenStudy (anonymous):

plug in -5 everywhere you see an x so you get -2(-5) + 4